# Local EGOP for Continuous Index Learning

### v2

@@ Abstract
1 Introduction
Some objectives of kernel engineering include the design of specialized kernels suited to particular data structures ((Odone et al., 2005; Kondor and Jebara, 2003; Joachims, 1998; Chapelle et al., 1999; Barla et al., 2002; Vishwanathan et al., 2010), etc.), statistical principles ((Genton, 2001; Schölkopf et al., 1997; Osborne, 2010), etc.), and problems of interest (Gong et al. (2024); Kokot and Luedtke (2025), etc.). In regression settings, a classical approach is local feature learning, in which kernels are augmented by differential information at points of interest ((Schmid and Mohr, 1997; Wallraven et al., 2003; Lowe, 1999)). Earlier nonparametric methods developed a similar framework, with datasets being recursively partitioned to improve the quality of local fits (Heise, 1971; Breiman and Meisel, 1976; Friedman, 1979).
Figure 1: Localizations from Local EGOP Learning (Algorithm 1) centered at each of the highlighted points (in red) trained on a grayscale image of a mandrill Bush (2021). Here $X$ is the pixel location and $Y$ the grayscale value of the image. For visualization purposes the procedure was stopped early to enforce large localizations, and the 125 highest weight pixels are highlighted at each point. On the right the image is magnified to the highlighted region boxed-off on the left.
A modern incarnation of kernel engineering is multi-index learning, particularly in the case of neural networks ((Mousavi-Hosseini et al., 2022; Boix-Adsera et al., 2023; Damian et al., 2023), etc.). This literature aims to show that the desirable properties of kernel engineering, such as data adaptivity and dimension reduction, are captured implicitly by certain machine learning models. Much work has been done in the single-index case, where the goal is to regress labels $Y$ on features $X$ when $f(x):=\mathbb{E}[Y|X=x]$ depends on $x$ solely through their evaluation in a fixed direction $v,$ via the quantity $v^{T}x$. It has been shown that two-layer neural networks not only learn this dependence, but also do so efficiently, leading to rapid increases in prediction quality ((Bietti et al., 2022; Abbe et al., 2024; Lee et al., 2024), etc.). Recent results indicate that these same benefits carry over to the multi-index setting for models trained via SGD (Damian et al., 2023; Arnaboldi et al., 2024).
In this paper, we generalize multi-index learning, introducing a new setting we call continuous-index learning (CIL). Rather than encoding the dependence of $x$ through the action of a low-rank matrix $V,$ $f(x)=g(Vx)$, we allow $V_{x}$ to be a continuously varying affine map, $f(x)=g(V_{x}x)$. Our window into this setting is via the supervised noisy manifold hypothesis, where we assume that the features $X$ are concentrated about a $d$-dimensional manifold $\mathcal{M}$, and $f(x)=g(\pi(x))$ for $\pi(x):=\operatorname{argmin}_{q\in\mathcal{M}}\|q-x\|$ the nearest point projection onto $\mathcal{M}$. For such data, the normal space $\mathcal{N}_{x}$ represents a continuously varying uninformative subspace, that is, for $\eta$ normal to $\mathcal{M}$ at $p:=\pi(x)$, $\|\eta\|$ sufficiently small, $f(x+\eta)=f(x)$. The tangent $\mathcal{T}_{x}=\mathcal{N}_{x}^{c}$ we refer to as the informative subspace. This feature structure is typical in manifold learning settings (Aamari and Levrard, 2018; Genovese et al., 2012; Kokot et al., 2025). The supervised noisy manifold hypothesis appears under a different name in the recent work Wu and Maggioni (2024), where the analysis is specialized to a 1-dimensional intrinsic manifold, a setting they call the “nonlinear single-variable model”.
@@ 1 Introduction
We approach this problem via kernel smoothing, $\hat{f}(x)=\sum_{i}w_{i}Y_{i}$, $w_{i}\propto k(\|x-X_{i}\|)$ for a rbf $k$. By introducing a Mahalanobis metrization $k(\|M^{1/2}(x-X_{i})\|)$, we seek to emulate the local multi-index structure of a continuous-index outcome. In particular, if $M$ can be selected to degenerate along the normal then the weights will decay primarily along the informative tangent direction, leading to estimation rates scaling with the intrinsic manifold dimension $d$. Without a priori knowledge of the manifold $\mathcal{M}$, we develop a recursive procedure we call Local EGOP Learning, which iteratively refines the metric $M$ by pooling coarse differential information estimated from the observed data.
Our method bears particular resemblance to “kernel steering”, which was developed in the image processing literature (Takeda et al., 2007) (see also the follow-up paper (Takeda et al., 2008)). The goal of kernel steering is to allow the kernel size and shape to change in a data-dependent way, adapting not only to sample location and density, but also to local features in the data. A special case of these data-adaptive kernels appears in conjunction with the popular bilateral filter in computer vision (Tomasi and Manduchi, 1998), (Elad, 2002). In Takeda et al. (2007), they propose an iterative procedure that estimates gradients about a point of interest. These are then used to “steer” the kernel locally, adopting the empirical covariance of the gradients $\hat{\mathcal{L}}$ near the point of interest as a Mahalanobis distance for subsequent estimation. Applying this method to the Gaussian kernel gives the steering kernel
$K_{h,\hat{\mathcal{L}}}(x_{i}-x^{*})\propto\exp{\left\{-\frac{(x_{i}-x^{*})^{T}\hat{\mathcal{L}}(x_{i}-x^{*})}{2h^{2}\mu^{2}}\right\}},$
which can then be used to estimate the function again and further refine the estimate of $\hat{\mathcal{L}}$. To the best of our knowledge, concrete guarantees have yet to be proven for this algorithm. This directional adaptation enables more effective denoising and image recovery, and is closely related to the expected gradient outer product (EGOP) $\mathcal{L}(\mu):=\mathbb{E}_{\mu}[\nabla f(X)\nabla f(X)^{T}]$ (Samarov, 1993; Hristache et al., 2001; Xia et al., 2002; Trivedi et al., 2014; Yuan et al., 2023). Besides kernel steering, local EGOP metrizations have appeared sporadically in different literatures, with common goals being to accelerate learning tasks and learn efficient dimension reductions (Trivedi et al., 2014; Wu et al., 2010; Mukherjee et al., 2010).
Recently, a similar procedure proposed in works such as Radhakrishnan et al. (2022, 2025); Zhu et al. (2025) emphasizes the importance of the Average Gradient Outerproduct (AGOP), which is the empirical covariance matrix of the estimated gradients $\hat{\mathcal{L}}(P_{n}):=\frac{1}{n}\sum_{i=1}^{n}\hat{\nabla}f\hat{\nabla}f^{T}$. These papers demonstrate, both empirically and theoretically, that this object is closely associated with the performance of simple neural networks, and that its inclusion in a kernel machine allows for estimation adaptive to the global regularity of functions of interest.
@@
However, $\mathcal{L}(\mu)$ may generically be full rank, making procedures such as the aforementioned subject to the curse of dimensionality. The supervised noisy manifold hypothesis is exactly such an example. Generically, $\cup_{x\in\mathcal{M}}\mathcal{T}_{x}=\mathbb{R}^{D}$, the full ambient space, hence globally all subspaces may be informative. Only locally will the EGOP degenerate, necessitating $\mu$ appropriately concentrate about the point of interest $x^{*}$. We study this through the lens of Gaussian localizations, $\mu=N(x^{*},\Sigma)$. The covariance $\Sigma$ and the Mahalanobis distance $M$ have inverted proclivities, as where $M$ seeks to degenerate along the normal, $\mathcal{L}(\mu)$ is most degenerate for $\Sigma$ aligned with the normal. This is no coincidence, as upon selecting $k$ Gaussian, smoothing with weights $\propto k(M^{1/2}(X_{i}-x^{*}))$ corresponds to an average with respect to $\mu=N(x^{*},M^{-1})$.
We formally link these objects in Section 3, developing a rigorous framework to assess the contribution of anisotropy to learning rates for kernel estimators. In sections 4 and 5, we demonstrate how optimizing the corresponding bias-variance trade-off naturally leads to the Local EGOP Algorithm, a recursive procedure reminiscent to the methods proposed in Takeda et al. (2007); Radhakrishnan et al. (2022). Finally, in section 6, we show that this method leads to adaptive regression guarantees for both noisy manifold and generic data. We support the theoretical results with numerical examples illustrating our claims in Section 7, and discuss further directions of research in Section 8.
2 Assumptions and Notation
@@
We refer to the population distribution of our feature vectors $X\in\mathbb{R}^{D}$ as $P$, and its empirical distribution as $P_{n}$. We assume $P$ has bounded $C^{1}$ density with respect to the Lebesgue measure on $\mathbb{R}^{D}$. When invoking the noisy manifold hypothesis, we denote by $\mathcal{M}$ the base manifold of dimension $d$, by $\pi$ the nearest point projection onto the manifold, and by $\mathcal{T}_{x}$ and $\mathcal{N}_{x}$ the tangent and normal spaces, respectively, at point $p:=\pi(x)\in\mathcal{M}$. For $\pi$ to be well-defined, we further assume that $X$ lies within the reach of $\mathcal{M}$ Federer (1959) almost surely. The labels $Y$ are assumed to be of the form $Y_{i}=f(X_{i})+\varepsilon$ for $\varepsilon_{i}$ iid, mean zero, and finite 4th moment. We assume that $f\in C^{4}$, bounded with uniformly bounded derivatives. We reserve $\mu$ to refer to a localization or target distribution of data appropriately concentrated about the point of interest $x^{*}$ with covariance $\Sigma$. We assume a uniform cap for the covariance, $\|\Sigma\|\leq\zeta<\infty$. We denote by $\langle\cdot,\cdot\rangle_{F}$ the Frobenius inner product. This work centers around the EGOP, $\mathcal{L}(\mu):=\int\nabla f(x)\nabla f(x)^{T}\,d\mu$, and its empirical counterpart the AGOP, $\hat{\mathcal{L}}(w):=\sum_{i}w_{i}\hat{\nabla}f\hat{\nabla}f^{T}$, for some estimator $\hat{\nabla}f$ of the gradient. We assume $g=\nabla f(x^{*})\neq 0$. By $k$, we denote a second-order kernel, satisfying typical assumptions in the nonparamtric statistics literature, see for example Tsybakov (2009)[Chapter 1]. In Section 3, we specialize results to the Gaussian kernel $k(\|x\|)\propto\exp(-\|x\|^{2})$, but we discuss the general setting in Appendix B.1.4.
3 Bias Control via EGOP

### v3

@@ Abstract
Machine Learning, ICML, EGOP, Continuous Index Learning
1 Problem and approach
Figure 1: Localizations from Local EGOP Learning (Algorithm 1) centered at each of the highlighted points (in red) trained on a grayscale image of a mandrill Bush (2021). Here the inputs $x$ are the pixel location and $f(x)$ is the grayscale value of the image. For visualization purposes the procedure was stopped early, and the 125 highest weight pixels are highlighted at each point. On the right the image is magnified to the highlighted region boxed-off on the left.
In a variety of contexts, including speech, images, and molecular dynamics, high-dimensional data often lie near a non-linear but low-dimensional manifold. Motivated by this phenomenon, we formulate the supervised noisy manifold hypothesis (SNMH), which posits that a target function $f(x)$ varies along, but not orthogonally to, such a manifold $\mathcal{M}$. We call the problem of estimating $f$ continuous index learning (CIL), analogous to the well-studied problem of multi-index learning (MIL), which assumes that $\mathcal{M}$ is a linear subspace.
Given observations $(x_{i},y_{i})$, $i\in[n]$, multi-index learning is the task of estimating a function $f(x):=\mathbb{E}[Y|X=x]$ that only depends on $x$ through composition with a low-rank matrix $V$, i.e. $f(x)=g(Vx).$ In CIL, we allow $V_{x}$ to be a continuously varying affine map, i.e. $f(x)=g(V_{x}x)$, and $\nabla V_{x}\eta=0$ for all $\eta$ such that $V_{x}(x+\eta)=V_{x}(x)$. Due to the SNMH, $f(x)=g(\pi(x))$ where $\pi(x):=\operatorname{argmin}_{q\in\mathcal{M}}\|q-x\|$ denotes the nearest point projection onto $\mathcal{M}$. For such data we call the space normal to $\mathcal{M}$, denoted $U_{\pi(x)}$, the uninformative subspace, as $f$ does not vary along these directions. We refer to the tangent directions, denoted $V_{x}=U_{\pi(x)}^{\perp}$, as the informative subspace.
@@ 1 Problem and approach
We approach estimating $f(x)$ via kernel smoothing, letting
$\hat{f}(x)=\sum_{i}w_{i}Y_{i},\;\text{with }w_{i}\propto k_{h}(x,X_{i})$ (1)
for a kernel $k$ and bandwidth $h$. For convenience, we use a Gaussian kernel for $k$, although modifications for the general setting are presented in Appendix B.1.4.
@@
We propose that the $k_{h}$ be anisotropic, emulating the local multi-index structure of $f$ through the Mahalanobis metrization $k_{h}(x,X_{i})=k(\|M^{1/2}(x-X_{i})\|/\sqrt{2h})$. Specifically, if $M$ can be selected to degenerate along the normal $U_{\pi(x)}$, then the weights $w_{i}$ will decay primarily along the informative tangent direction $V_{x}$, leading to estimation rates scaling with the dimension $d$ of the intrinsic manifold $\mathcal{M}$.
We then develop a recursive procedure we call Local EGOP Learning, which iteratively refines the metric $M$ by pooling coarse differential information estimated from the observed data around a target point $x^{*},$ without a priori knowledge of the manifold $\mathcal{M}$.
Our method draws inspiration from recent works such as Radhakrishnan et al. (2022); Beaglehole et al. (2023); Radhakrishnan et al. (2025); Zhu et al. (2025); more background can be found in Section 2. Central in this line of research is the expected gradient outer product (EGOP), $\mathcal{L}(\mu):=\int\nabla f\nabla f^{T}\,d\mu$. Its empirical counterpart is the average gradient outer product (AGOP), $\hat{\mathcal{L}}(P_{n}):=\mathcal{L}_{\hat{f}}(P_{n}):=\frac{1}{n}\sum_{i=1}^{n}\nabla\hat{f}(x_{i})\nabla\hat{f}^{T}(x_{i})$, with $P_{n}$ the empirical distribution, $\hat{f}$ an estimator of $f$. In this paper, we focus on the problem of point estimation, where $f$ is estimated at a specified query point $x^{*}$. Our key insight is to iteratively localize the EGOP (or AGOP) around the point of interest $x^{*}$. At a given iteration $t$, we set $w_{i}^{(t)}\propto k(\|\hat{M}_{t}^{1/2}(x^{*}-X_{i})\|/\sqrt{2h_{t+1}})$, use $\hat{f}$ given by (1), and compute
$\hat{\mathcal{L}}_{t}:=\hat{\mathcal{L}}\left(\sum_{i=1}^{n}w_{i}^{(t)}\delta_{x_{i}}\right)\;=\sum_{i=1}^{n}w_{i}^{(t)}\nabla\hat{f}(x_{i})\nabla\hat{f}(x_{i})^{T}.$
That is, rather than averaging the gradient outer products over the full dataset, we weigh them with respect to the same kernel we utilize for regression. In the subsequent iteration, we update the metric $\hat{M}_{t+1}$ using $\hat{\mathcal{L}}_{t}$, thus steering the kernel with this differential information. This recursion is at the core of Algorithm 1 described in Section 3,
In Section 4, we analyze this recurrence through its continuum counterparts, $\mu_{t+1}:=N(x^{*},h_{t+1}M_{t}^{-1})$, $M_{t}:=\beta\mathcal{L}(\mu_{t})+(1-\beta)\mathcal{L}(\mu_{t-1})$. In Section 5, we numerically validate these results, and compare the performance of our proposed estimator to various neural network architectures, while Section 6 indicates avenues for future work. Our main contributions are:
  1. 1.
    Introducing the setting of continuous index learning, generalizing multi-index learning to latent manifold structures.
  2. 2.
    Developing the Local EGOP Learning Algorithm and the localization of the EGOP.
  3. 3.
    Analyzing the recurrence and the resulting intrinsic dimensional learning rate of Local EGOP Learning under the SNMH (Sections 4.3 and 4.4).
  4. 4.
    Establishing a novel approach for the study of recursive kernel learning such as Radhakrishnan et al. (2022, 2025); Zhu et al. (2025).
  5. 5.
    Clarifying the role of the EGOP in the minimization of the mean squared error via a new functional framework.
@@
2 Background and Related Work
Some objectives of kernel engineering include the design of specialized kernels suited to particular data structures (Odone et al. (2005); Kondor and Jebara (2003); Joachims (1998); Chapelle et al. (1999); Barla et al. (2002); Vishwanathan et al. (2010), etc.), statistical principles (Genton (2001); Schölkopf et al. (1997); Osborne (2010), etc.), and problems of interest (Gong et al. (2024); Kokot and Luedtke (2025), etc.). In regression settings, a classical approach is local feature learning, in which kernels are augmented by differential information at points of interest (Schmid and Mohr (1997); Lowe (1999); Wallraven et al. (2003)). Earlier nonparametric methods developed a similar framework, with datasets being recursively partitioned to improve the quality of local fits (Heise, 1971; Breiman and Meisel, 1976; Friedman, 1979).
A modern incarnation of kernel engineering is multi-index learning, particularly in the case of neural networks (Mousavi-Hosseini et al. (2022); Boix-Adsera et al. (2023); Damian et al. (2023), etc.). This literature aims to show that the desirable properties of kernel engineering, such as data adaptivity and dimension reduction, are captured implicitly by certain machine learning models. Much work has been done in the single-index case, where the goal is to regress labels $Y$ on features $X$ when $f$ depends on $x$ solely through its evaluation in a fixed direction $v,$ via the quantity $v^{T}x$. It has been shown that two-layer neural networks not only learn this dependence, but do so efficiently, leading to rapid increases in prediction quality (Bietti et al. (2022); Abbe et al. (2024); Lee et al. (2024), etc.). Recent results indicate that these same benefits carry over to the multi-index setting (Damian et al., 2023; Arnaboldi et al., 2024).
The SNMH feature structure is typical in manifold learning (Aamari and Levrard, 2018; Genovese et al., 2012; Kokot et al., 2025). It also appears under a different name (the “nonlinear single-variable model”) in the recent work (Wu and Maggioni, 2024), where the analysis is specialized to a 1-dimensional intrinsic manifold. They relate this problem to the more general settings of compositional learning (Juditsky et al. (2009); Shen et al. (2021), etc.) and non-linear dimension reduction (Yeh et al. (2008); Lee et al. (2013), etc.). In this literature, many recent methods propose similarly motivated optimization schemes that target related Poincaré inequalities (Bigoni et al., 2022; Verdière et al., 2025; Nouy and Pasco, 2025).
Our method bears a particularly strong resemblance to “kernel steering” as proposed in the image processing literature (Takeda et al., 2007, 2008). The EGOP has been motivated in many statistical applications (Samarov, 1993; Hristache et al., 2001; Xia et al., 2002; Trivedi et al., 2014; Yuan et al., 2023). A precursor in a supervised noiseless manifold setting is Aswani et al. (2011); in this paper, intrinsic rates are achieved with an isotropic kernel and sparse regression, where the regression coefficients are penalized differently in the normal and tangent space to $\mathcal{M}$. Unlike our work and much of the literature, the method is extended to the regime where the number of regressors tends to infinity. The methods are not adaptive, an outer procedure must estimate the local dimension $d$.
Besides kernel steering, EGOP metrizations have appeared in many different literatures, with common goals being to accelerate learning tasks and learn efficient dimension reductions (Trivedi et al., 2014; Wu et al., 2010; Mukherjee et al., 2010).
3 The Local EGOP Algorithm
In this section, we introduce the Local EGOP Learning algorithm, detailed in Algorithm 1.
At iteration $t$, the algorithm uses the Gaussian kernel with Mahalanobis metric $M_{t}$ (step 3) to estimate the local density $\mu_{t}$ about $x^{*}$. Using this weighting, $\hat{f}$ and its gradients are computed (steps 5–7), which in turn are used to calculate the AGOP, $L_{t}$. This is then incorporated into the new metric $M_{t+1}=[\beta L_{t}+(1-\beta)L_{t-1}]$, $\beta\in(0,1)$.
Algorithm 1 Local EGOP Learning
0: Data $\{(X_{j},Y_{j})\}_{i=1}^{n}$, target $x^{*}$, subsample size $m$, iterations $T$, bandwidths $\{h_{t}\}$, $\beta>0$, scale $\alpha>0$.
0: Best estimate $\hat{f}(x^{*})$.
[…]
2:for $t=0$ to T1T-1 do
3:  $\triangleright$ Compute weights at $x^{*}$
4:  $w_{i}\propto\exp\!\bigl(-(x_{i}-x^{*})^{\top}M_{t}(x_{i}-x^{*})\bigr)$ $i=1,\dots,n$
5:  Normalize $w$, subsample $S\subset\{1,\dots,n\},|S|=m$ with probs $w$
6:  $\triangleright$ Leave-One-Out Local regressions at each $i\in S$
7:  for $i\in S$ do
8:   Fit weighted local linear regression centered at $x_{i}$ (weights from metric $M_{t}$), excluding $(x_{i},y_{i})$
9:   Obtain $\nabla\hat{f}[i]$ and $\hat{f}[i]$
10:  end for
11:  $\triangleright$ Aggregate statistics
12:  $L_{t}\leftarrow\sum_{i\in S}w_{i}\,\nabla\hat{f}[i]\nabla\hat{f}[i]^{\top}$
13:  $\widehat{\mathrm{MSE}}_{t}\leftarrow\sum_{i\in S}w_{i}\,(y_{i}-\hat{f}[i])^{2}$
14:  if $\widehat{\mathrm{MSE}}_{t}<\text{bestMSE}$ then
15:   Estimate $\hat{f}(x^{*})$ via local regression centered at $x^{*}$
16:   $\text{bestMSE}\leftarrow\widehat{\mathrm{MSE}}_{t}$,  $\hat{f}_{\text{best}}(x^{*})\leftarrow\hat{f}(x^{*})$
17:  end if
18:  $\triangleright$ Metric update: set $\beta=1$ if $i=0$, $\beta<1$ otherwise
19:  $M_{t+1}\leftarrow\bigl(\beta L_{t}+(1-\beta)L_{t-1}\bigr)$
20:  $M_{t+1}\leftarrow M_{t+1}/[t_{t+1}\operatorname{tr}(M_{t+1})]$
21:end for
22:return $\hat{f}_{\text{best}}(x)$
In practice, one might choose hyper-parameters $\beta,T$, as well as the schedule $\{h_{t}\}$ via cross-validation. For computational reasons, we normalize $M_{t}$ by its trace (step 20) when applying the scaling factor $h_{t+1}$. This has little impact asymptotically, as the trace converges to the leading gradient eigenvalue of the EGOP, making this normalization effectively equivalent to a constant scaling. However, it serves a practical benefit by stabilizing initial iterates, setting the scale of the metric to the predefined rate $1/h_{t}$. Note that while the optimal number of iterations is dimension-dependent, as argued in Theorems 1 and 3, our algorithm incorporates a Leave-One-Out style validation loop (step 10) for adaptive stopping. Further theoretical motivation for the algorithm is presented in Section 4.3.
The time complexity of Algorithm 1 is $O(T((m+2)nD^{2}+(m+1)D^{3}))$, with memory overhead $O(nD^{2}+m^{2}D^{2}+mD^{3})$.
4 Theoretical analysis
We now study Algorithm 1 through its continuum $t$ counterparts, described by $\mu_{t+1}:=N(x^{*},h_{t+1}M_{t}^{-1})$ and $M_{t}:=\beta\mathcal{L}(\mu_{t})+(1-\beta)\mathcal{L}(\mu_{t-1})$. Setting $\Sigma_{t}:=\Sigma(\mu_{t}):=\int(x-x^{*})(x-x^{*})^{T}\,d\mu_{t}$, this reduces to the matrix recurrence
$\Sigma_{t+1}=h_{t+1}[\beta\mathcal{L}(\mu_{t})+(1-\beta)\mathcal{L}(\mu_{t-1})]^{-1}.$ (2)
In Section 4.2 we propose a framework for the analysis of the resulting estimator. We introduce the EGOP form $W(\mu_{t}):=\operatorname{tr}(\mathcal{L}(\mu_{t})\Sigma(\mu_{t}))$, relating it to the bias of the $M_{t}$ metrized kernel smoother via a Poincaré inequality. Building on this, in Section 4.3, we demonstrate how this recursive kernel adaptation can be naturally framed as a greedy MSE reduction strategy. In Section 4.4, our analysis yields an intrinsic learning rate for supervised noisy manifold data, and a significant efficiency gain otherwise.
4.1 Assumptions and Notation
@@
We refer to the population distribution of our feature vectors $X\in\mathbb{R}^{D}$ as $P$, and its empirical distribution as $P_{n}$. We assume $P$ has bounded $C^{1}$ density with respect to the Lebesgue measure on $\mathbb{R}^{D}$. When invoking the noisy manifold hypothesis, we denote by $\mathcal{M}$ the smooth base manifold of dimension $d$, by $\pi$ the nearest point projection onto the manifold, and by $\mathcal{T}_{x}$ and $\mathcal{N}_{x}$ the tangent and normal spaces $U_{\pi(x)}$ and $V_{x}$, respectively, at point $p:=\pi(x)\in\mathcal{M}$. For $\pi$ to be well-defined, we further assume that $X$ lies within the reach of $\mathcal{M}$ (Federer, 1959) almost surely. The labels $Y$ are assumed to be of the form $Y_{i}=f(X_{i})+\varepsilon_{i}$ for $\varepsilon_{i}$ iid, mean zero, and finite 4th moment. We assume that $f\in C^{4}$ and is bounded with uniformly bounded derivatives. We reserve $\mu$ to refer to a localization or target distribution of data appropriately concentrated about the point of interest $x^{*}$. We denote by $\Sigma$ the covariance of $\mu$, and we enforce $\|\Sigma\|\leq\zeta<\infty$. We denote by $\langle\cdot,\cdot\rangle_{F}$ the Frobenius inner product. We assume $g=\nabla f(x^{*})\neq 0$. We denote by $k$ a second-order kernel that satisfies typical assumptions in the nonparamtric statistics literature, as in, for example, (Tsybakov, 2009)[Chapter 1].
4.2 Bias Control via EGOP